# Flow Snapshot Neurons in Action: Deep Neural Networks Generalize to Biological Motion Perception

### v1

@@ 3.2 Video Action Recognition Datasets and Baselines in Computer Vision
To evaluate the ability of all AI models to recognize a wider range of actions in natural RGB videos, we include two standard video action recognition datasets in computer vision. NTU RGB+D 60 [80] comprises 56,880 video clips featuring 60 action classes. NW-UCLA [92] contains 1,493 videos featuring 10 action classes. Both datasets are split randomly, with a 7:3 ratio for the training and test sets. All AI models are trained on RGB videos and tested on Joint videos of these datasets curated with Alphapose [25] described above. We compute the top-1 action recognition accuracy of all AI models. In addition to our MP model, we include six competitive baselines below for comparison: ResNet3D [35] adapts the architecture of ResNet [36] by substituting the 2D convolutions with the 3D convolutions. I3D [11] extends a pre-trained 2D-CNN on static images to 3D by duplicating 2D filters along temporal dimensions and fine-tuning on videos. X3D [27] expands 2D-CNN progressively along spatial and temporal dimensions with the architecture search. R(2+1)D [88] factories the 3D convolution kernels into spatial and temporal kernels. SlowFast [28] is a two-stream architecture that processes temporal and spatial cues separately. MViT [24] is a transformer-based model that expands feature map sizes along the temporal dimension while reducing them along the spatial dimension.
@@ 4.1 Our model achieves human-level performance without task-specific retraining
Figure 4: Temporal orders and resolutions matter in generalization performance on RGB and Joint videos. Stimuli encompass RGB and Joint (J) videos. Short forms include R (reversal), S (shuffle), F (frames), and P (points) in Sec. 3.1. Error bars indicate the standard error.
Figure 5: Our model demonstrates human-like robustness to minimal visual information. Top-1 action recognition accuracy is a function of the number of joints (P) in Joint (J) videos. Results from RGB test videos are at the leftmost for comparison.
We explore five key properties of motion perception and their influence on video action recognition tasks between humans and our model (MP). Detailed comparisons between humans and all AI models across all BMP tasks are provided in the Tab. S1.
@@
Moreover, we observed that shuffled (S) or reversed (R) temporal orders significantly impair recognition performance in both humans and MP. This effect is more pronounced in Joint videos (J-6P-R and J-6P-S) compared to RGB videos (RGB-R and RGB-S). The minimal motion information available in Joint videos makes them particularly susceptible to disruptions in temporal orders. The same behavioral patterns are also captured by MP. Interestingly, shuffling has a lesser impact on human performance compared to reversing RGB videos (RGB-R versus RGB-S). However, this pattern is reversed in Joint videos (J-6P-R versus J-6P-S). When considering actions such as sitting down versus standing up, reversing orders may alter the temporal context in RGB videos. The effect of temporal continuity outweighs the effect of temporal context in Joint videos.
@@
Minimal amount of visual information is sufficient for recognition. In Fig. 5, both humans and all AI models exhibit comparable performances in RGB videos (RGB). There is a slight decrease in accuracy when transitioning from RGB to J-6P for both humans and MP. Interestingly, the accuracy drop from J-6P to J-5P is indistinguishable for humans. Similarly, there was a wide range of number of points that led to robust action recognition for MP (J-26P to J-5P). These observations suggest that a minimal visual representation in J-5P is sufficient for action recognition. Conversely, traditional AI models show a significant performance decline moving from J-26P to J-5P. These findings suggest that existing AI models struggle with reduced visual information.
@@
Camera views have minimal impacts on recognition. Neuroscience studies  [38, 89, 41, 61] provide evidence that both viewpoint-dependent and viewpoint-invariant representations for action recognition are encoded in primate brains. In Fig.8, MP exhibits a slightly higher accuracy on frontal and $45^{\circ}$ views compared to the $90^{\circ}$ (profile) view by 2.5%, which mirrors human performance patterns in [41]. However, we note that human behaviors in our study differ slightly from  [41] (see Sec.B). Moreover, MP significantly outperforms the second-best model, MViT, by 18%. This suggests that MP not only captures human-like viewpoint-dependent representations but also maintains superior accuracy among AI models across different camera views.
@@ 4.2 Comparisons among AI models in BMP tasks and standard computer vision datasets
MP significantly outperforms all the competitive baselines in BMP tasks. In Fig. 7, we reported the correlation between each AI model and human performance across five fundamental properties of motion perception. We averaged the accuracies within the same property. It is observed that all AI models demonstrated lower accuracy than humans. However, among all AI models, ours not only outperforms the rest but also shows the highest correlation with human performance. For example, despite explicitly modelling both motion and visual information in two separate streams, the SlowFast model still falls short in temporal orders. Similarly, MViT, which utilizes transformer architectures, fails to generalize to various temporal resolutions. See Tab. S1 for detailed comparisons.
Figure 6: Both humans and our model exhibit the ability to recognize actions in sequential position actor (SP) videos without local image motions. Performance varies depending on the persistence of visual information, with stimuli having 4 and 8 points (P) of the actors. See Sec. 3.1 for task details.
Figure 7: Our model significantly surpasses all existing models on both Joint and SP videos. A correlation plot between model and human performances is provided, with accuracy averaged over all conditions within each property on the same type of stimulus (Sec. 3.1). The black dash diagonal serves as a reference.
@@
MP outperforms all the competitive baselines on Joint videos from two standard computer vision datasets. In Tab. 1, MP, relying solely on patch-level optical flows as inputs, performs above chance. This suggests that motion information alone is sufficient for accurate action recognition. Moreover, MP performs better than existing models on Joint videos in these datasets, highlighting that MP learns to capture better motion representations from RGB videos during training. The MP model explicitly uses patch-level optical flow to extract temporal information for classification, whereas the baselines do not. Thus, here we introduce augmented baselines where they are trained on pixel-level optical flows. Specifically, we use GMFlow [97, 98] to extract the pixel-level optical flow from RGB videos in the BMP dataset, and then convert them into three-channel videos based on the Middlebury colour code [5]. Taking these optical flow videos in the training set as the input to baselines, the performances on optical flow videos in the testing set and J-6P are shown in Tab. 2. It is evident that all baselines achieve high accuracy on optical flow videos in the testing set, indicating that they effectively learn to recognize actions from optical flow inputs. However, their best performances on J-6P are significantly worse compared to the MP model (30.83% vs 69.00% in Tab. S1), suggesting that our MP learns more generalizable motion representations from patch-level optical flows.
[…]
Table 2: Results of baselines with optical flow as input. Top-1 accuracy is reported on optical flow videos (OF) and Joint videos with 6 points (J-6P). Best is in bold.
@@ 4.3 Ablation studies reveal key components in our model
SS FSN Slots $\hat{Z}$ Thres. $\gamma$ T Emb. $L_{slot}$ T Aug. Top-1 Acc. (%)
@@
Table 3: Ablation reveals critical components in our model. Top-1 accuracy is reported on RGB videos, Joint videos with 6 points (J-6P), and SP videos with 8 points and a lifetime of 1 (SP-8P-1LT). From left to right, the ablated components are: the pathway with Motion Invariance Neurons (MIN), the pathway with Flow Snapshot Neurons (FSN), the single-scale branch with FSN $\hat{M}_{1}$, FSN $\hat{Z}$, threshold $\gamma$ (Sec. 2.1), the time embedding (Sec. 2.4), the loss term $L_{slot}$ (Sec. 2.2) and data augmentation by randomly shuffling and reversing training frames within the same video. Best is in bold.
@@
To investigate how different components of MP contribute to the generalization performance on Joint and SP videos, we conducted ablation studies (Tab. 3 and Tab. 4). Removing the pathway involving motion invariance neurons (A1, Tab. 3) leads to a drop in accuracy. Similarly, removing the pathway involving flow snapshot neurons also leads to the accuracy drop (A2). Both ablation studies highlight the importance of two pathways for action recognition.
@@
Instead of multi-scale feature fusion, we utilize a single-scale pathway of flow snapshot neurons (A3). The decrease of 10.57% in accuracy on J-6P videos implies that multi-scale flow snapshot neurons capture both fine-grained and high-level temporal patterns, essential for action recognition. In A4, the model takes patch-level optical flows as inputs and directly utilizes them for feature fusion and action classification without flow snapshot neurons. There is a significant drop of 14.33% in accuracy, implying that flow snapshot neurons capture meaningful prototypical flow patterns, and their activations are critical for recognition. Threshold $\gamma$ controls the level of noise or errors in small motions. As expected, removing $\gamma$ leads to impaired recognition accuracy (A5). As illustrated in Sec. 2.4, time embeddings are appended to the activations of flow snapshot neurons. Aligning with [59], time embeddings provide useful temporal context (A6). The loss $L_{slot}$ on contrastive walks encourages the slots to capture diverse prototypical flow patterns. Consistent with [94], removal of $L_{slot}$ (A7) leads to degraded generalisation performance, indicating the importance of contrastive walk loss between slots and optical flows.
@@
Data augmentation, a common technique to enhance AI model generalization [81], includes temporal augmentations like random shuffling and reversal in A8. Surprisingly, MP’s generalization performance is impaired, indicating that randomizing frame orders during training disrupts motion perception and action understanding. We also present ablation results in RGB videos, with similar conclusions drawn, albeit to a lesser extent due to increased visual information and reduced reliance on motion.
4.4 Ablation Study of Hyper-parameters

### v2

@@ 3.2 Video Action Recognition Datasets and Baselines in Computer Vision
To evaluate the ability of all AI models to recognize a wider range of actions in natural RGB videos, we include two standard video action recognition datasets in computer vision. NTU RGB+D 60 [87] comprises 56,880 video clips featuring 60 action classes. NW-UCLA [100] contains 1,493 videos featuring 10 action classes. Both datasets are split randomly, with a 7:3 ratio for the training and test sets. All AI models are trained on RGB videos and tested on Joint videos of these datasets curated with Alphapose [26] described above. We compute the top-1 action recognition accuracy of all AI models. In addition to our MP model, we include six competitive baselines below for comparison: ResNet3D [38] adapts the architecture of ResNet [39] by substituting the 2D convolutions with the 3D convolutions. I3D [11] extends a pre-trained 2D-CNN on static images to 3D by duplicating 2D filters along temporal dimensions and fine-tuning on videos. X3D [28] expands 2D-CNN progressively along spatial and temporal dimensions with the architecture search. R(2+1)D [96] factories the 3D convolution kernels into spatial and temporal kernels. SlowFast [29] is a two-stream architecture that processes temporal and spatial cues separately. MViT [25] is a transformer-based model that expands feature map sizes along the temporal dimension while reducing them along the spatial dimension. Furthermore, we include three more competitive baselines (E2-S-X3D [42], VideoMAE [94], TwoStream-CNN [90]) and present results in Appendix, Sec. E. The findings remain consistent with the baselines above. All baselines except for TwoStream-CNN are pre-trained on Kinetics 400 [47]. TwoStream-CNN is pre-trained on ImageNet [19]. As an upper bound, we also train the MP directly on Joint or SP videos and the results are in Appendix, Sec. E.
@@ 4.1 Our model achieves human-level performance without task-specific retraining
Figure 3: Temporal orders and resolutions matter in generalization performance on RGB and Joint videos. Stimuli encompass RGB and Joint (J) videos. Short forms include R (reversal), S (shuffle), F (frames), and P (points) in Sec. 3.1. Error bars indicate the standard error of the top-1 accuracy across different action classes.
Figure 4: Our model demonstrates human-like robustness under reduced visual information. Top-1 action recognition accuracy is a function of the number of points (P) in Joint (J) videos. Results from RGB test videos are at the leftmost. The colored shaded region represents the standard error across all action classes.
We explore five key properties of motion perception and their influence on video action recognition tasks between humans and our model (MP). Detailed comparisons between humans and all AI models across all BMP tasks are provided in the Appendix, Sec. L.
@@
Moreover, we observed that shuffled (S) or reversed (R) temporal orders significantly impair recognition performance in both humans and MP. This effect is more pronounced in Joint videos (J-6P-R and J-6P-S) compared to RGB videos (RGB-R and RGB-S). The minimal motion information available in Joint videos makes them particularly susceptible to disruptions in temporal orders. The same behavioral patterns are also captured by MP. Interestingly, shuffling has a lesser impact on human performance compared to reversing RGB videos (RGB-R versus RGB-S, p-value < 0.05). However, this pattern is reversed in Joint videos (J-6P-R versus J-6P-S, p-value < 0.05). When considering actions such as sitting down versus standing up, reversing orders may alter the temporal context in RGB videos. The effect of temporal continuity outweighs the effect of temporal context in Joint videos.
@@
Minimal amount of visual information is sufficient for recognition. In Fig. 4, both humans and all AI models exhibit comparable performances in RGB videos (RGB). Interestingly, the accuracy drop from J-6P to J-5P is indistinguishable for humans (p-value > 0.05). Similarly, there was a wide range of the number of points that led to robust action recognition for MP (J-26P to J-5P). However, its absolute accuracy is much lower than humans. Therefore, we introduce an enhanced version of the MP model (En-MP) by extending it across multiple DINO blocks, in addition to the last DINO block. Detailed implementation of the En-MP is provided in Appendix, Sec. F. Surprisingly, the En-MP outperforms our original MP significantly and performs competitively well as humans. These observations suggest that a minimal visual representation in J-5P is sufficient for action recognition on humans and our En-MP. Conversely, traditional AI models show a significant performance decline moving from J-26P to J-5P. These findings suggest that existing AI models struggle with reduced visual information.
@@
Camera views have minimal impacts on recognition. Neuroscience studies  [43, 97, 46, 68] provide evidence that both viewpoint-dependent and viewpoint-invariant representations for action recognition are encoded in primate brains. In Fig.7, MP exhibits a slightly higher accuracy on frontal and $45^{\circ}$ views compared to the $90^{\circ}$ (profile) view by 2.5%, which mirrors human performance patterns in [46]. However, we note that human behaviors in our study differ slightly from  [46] (see Appendix, Sec.G). Moreover, MP significantly outperforms the second-best model, MViT, by 18%. This suggests that MP not only captures human-like viewpoint-dependent representations but also maintains superior accuracy among AI models across different camera views.
@@ 4.2 Comparisons among AI models in BMP tasks and standard computer vision datasets
MP aligns with human behaviors more closely than all the competitive baselines in BMP tasks. In Fig. 6, we reported the correlation coefficients between each AI model and human performance across all conditions in the BMP dataset. The results demonstrate that our MP and En-MP exhibit a significantly higher correlation with human performance compared to the baselines, indicating that our MP and En-MP align more closely with human performance. In addition to the correlation coefficients, we also present the error pattern consistency [31] between models and humans (Appendix, Sec. H, Appendix, Fig. S4). Results suggest that our MP achieves the highest error pattern consistency with humans at the trial level.
MP significantly outperforms all the competitive baselines in BMP tasks. In Appendix, Fig. S3, we present the absolute accuracy of all models and humans. The slopes near 1 in our MP model indicate that it performs on par with humans and surpasses all baselines across the five BMP dataset properties. Despite explicitly modelling motion and visual information in separate streams, the SlowFast model struggles with temporal order. Likewise, MViT, which leverages transformer architectures, fails to generalize across different temporal resolutions.
Figure 5: Both humans and our model can recognize actions in SP videos without local motions. Performance varies depending on the persistence of visual information, with stimuli having 4 and 8 points (P) of the actors (Sec. 3.1).
Figure 6: Our MP model shows a significantly stronger correlation with human performance compared to all baselines. The correlation between the model and human performance across all BMP conditions is presented.
@@
MP outperforms all the competitive baselines on Joint videos from two standard computer vision datasets. In Tab. 1, MP, relying solely on patch-level optical flows as inputs, performs above chance. This suggests that motion information alone is sufficient for accurate action recognition. Moreover, MP performs better than existing models on Joint videos in these datasets, highlighting that MP learns to capture better motion representations from RGB videos during training. In Appendix, Sec. I, we provide results and discussions when we explicitly feed pixel-level optical flows as inputs to these baselines during training and test these models on the Joint videos. Our results show that MP still outperforms these models, suggesting that MP learns generalizable motion representations beyond optical flow patterns. In Appendix, Sec. J, we also present visualizations of patch-level optical flow examples, demonstrating its ability to semantically segment the movements of a person.
@@ 4.3 Ablation studies reveal key components in our model
SS FSN Slots $\hat{Z}$ Thres. $\gamma$ T Aug. Top-1 Acc. (%)
@@
Table 2: Ablation reveals critical components in our model. Top-1 accuracy is reported on RGB videos, Joint videos with 6 points (J-6P), and SP videos with 8 points and a lifetime of 1 (SP-8P-1LT). From left to right, the ablated components are: the pathway with Motion Invariance Neurons (MIN), the pathway with Flow Snapshot Neurons (FSN), the single-scale branch with FSN $\hat{M}_{1}$, Slots $\hat{Z}$, threshold $\gamma$ (Sec. 2.1), and data augmentation by randomly shuffling and reversing training frames within the same video. Best is in bold.
@@
To investigate how different components of MP contribute to the generalization performance on Joint and SP videos, we conducted ablation studies (Tab. 2, Appendix, Tab. S4 and Appendix, Tab. S5). Removing the pathway involving motion invariance neurons leads to a drop in accuracy when video frames are shuffled or reversed (A1, Tab. 2 and Appendix, Fig. S3). For example, MP outperforms its ablated model without MIN in the following experiments: 62.32% vs 49.58% in RGB-R; 61.34% vs 38.03% in RGB-S, 38.69% vs 36.03% in J-6P-R, and 32.65% vs 25.28% in J-6P-S. Similarly, removing the pathway involving flow snapshot neurons also leads to a significant accuracy drop (A2). Both ablation studies highlight the importance of two pathways for action recognition.
@@
Instead of multi-scale feature fusion, we utilize a single-scale pathway of flow snapshot neurons (A3). The decrease of 10.57% in accuracy on J-6P videos implies that multi-scale flow snapshot neurons capture both fine-grained and high-level temporal patterns, essential for action recognition. In A4, the model takes patch-level optical flows as inputs and directly utilizes them for feature fusion and action classification without flow snapshot neurons. There is a significant drop of 14.33% in accuracy, implying that flow snapshot neurons capture meaningful prototypical flow patterns, and their activations are critical for recognition. Threshold $\gamma$ controls the level of noise or errors in small motions. As expected, removing $\gamma$ leads to impaired recognition accuracy (A5).
@@
Data augmentation is a common technique to enhance AI model generalization [88, 110, 17, 40, 114, 49]. It typically includes temporal augmentations like random shuffling and reversal in A6. Surprisingly, MP’s generalization performance is impaired, indicating that randomizing frame orders during training disrupts motion perception and action understanding. We also present ablation results in RGB videos, with similar conclusions drawn, albeit to a lesser extent due to increased visual information and reduced reliance on motion.
In addition to the main ablation results discussed here, we also provide the extra ablation studies and model analysis in the Appendix. Specifically, we include the following studies: 1. the removal of the time embedding in our MP (Sec. 2.4, Appendix, Tab. S5); 2. the removal of the loss term $L_{slot}$ in our MP (Sec. 2.2, Appendix, Tab. S5); 3. the choice of the reference frame for calculating path-level optical flows in our MP (Sec. 2.1, Appendix, Tab. S5); 4. the pixel-level optical flows downscaled to the size of the patch-level optical flows in our MP (Appendix, Sec. K); 5. the replacement of the feature extractor DINO with ResNet in our MP (Appendix, Sec. K); 6. the comparison of the number of trainable parameters among all the models (Appendix, Sec. K); 7. the analysis of key frames predicted by our MP on an example video in Appendix, Sec. K. All these studies emphasized the importance of our model designs and demonstrated the generalization ability of our model.
5 Discussion
We introduce Motion Perceiver (MP) as a generalization model trained on natural RGB videos, capable of perceiving and identifying actions of living beings solely from their minimal motion patterns on point-light displays, even without prior training on such stimuli. Within MP, flow snapshot neurons learn prototypical flow patterns through competitive binding, while motion-invariant neurons ensure robustness to variations in temporal orders. The fused activations from both neural populations enable action recognition.
To evaluate the generalization capabilities of all AI models, we curated a dataset comprising  63k stimuli across 24 BMP conditions using point-light displays inspired by neuroscience. Psychophysics experiments on this dataset were conducted, providing human behavioral data for comparison with computational models. While AI models can surpass human performance in numerous tasks, current AI models for action recognition still fall short of human capabilities in many BMP conditions. By focusing solely on motion information, MP achieves superior generalization performance among all AI models and demonstrates a strong alignment with human behavioral responses. All baselines are pre-trained on large-scale video datasets, whereas our MP uses feature extractors pre-trained on naturalistic images. Remarkably, despite lacking video pre-training, our MP outperforms all baselines.
Our work takes an initial step toward bridging artificial and biological intelligence in BMP. First, it raises intriguing questions in neuroscience, such as the neural basis of motion-invariant neurons, which are crucial when video frames are shuffled or reversed. Bio-inspired architectures can help test specific neuroscience hypotheses, while insights from neuroscience can, in turn, guide the design of more advanced AI systems. Second, our work paves the way for real-world applications requiring robust motion recognition, such as in low-light conditions where visual information is limited.
The 10 action classes in our BMP dataset were selected for their rich temporal information. We observe significant variations in action recognition accuracy across various action classes for both human participants and AI models. In the future, the BMP dataset could be expanded to include more complex actions. Both neuroscience and computer vision use various biological motion stimuli; here, we focus on point-light displays, but future work could explore other visual stimuli, such as motion patterns in noisy backgrounds. While our model shows promising generalization on BMP stimuli, it does not account for attention or top-down influences. Moreover, since it only uses patch-level optical flows, integrating information from the ventral (form perception) and dorsal (motion perception) pathways remains an open challenge. Further discussion on the social impact of our work can be found in Appendix, Sec. M.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-025), and its NRFF award NRF-NRFF15-2023-0001. We also acknowledge Mengmi Zhang’s Startup Grant from Agency for Science, Technology, and Research (A*STAR), Startup Grant from Nanyang Technological University, and Early Career Investigatorship from Center for Frontier AI Research (CFAR), A*STAR.
Appendix A Biological Motion Perception (BMP) Dataset
A.1 Example Videos in the BMP Dataset
We provide one example RGB video and its corresponding BMP videos under different conditions where the subject is performing a sit-down action. The naming convention of these BMP videos follows the same as in Sec. 3.1. See the example video at the link.
A.2 Action Classes in the BMP Dataset
There are 10 action classes in the BMP dataset: pick up, throw, sit down, stand up, kick something, jump up, point to something, nod head/bow, falling down and arm circles.
A.3 More Implementation Details of the BMP Dataset
In Sec. 3.1, we vary the amount of visual information based on the number of point lights on Joint videos. To generate Joint videos, Alphapose [26] is used as the tool to detect the joints of a human body from RGB videos. After that, we use these detected joints to generate BMP stimulus on point-light displays. The specific positions of light points correspond to certain joint locations:
  • J-5P: 1 point each on the nose, 1 point each on one hand and one ankle.
We also looked into the invariance property to camera views in Sec. 3.1. The video clips in J-6P are categorized based on the viewpoints in the videos: the frontal view, $45\degree$ view and $90\degree$ view, which are respectively labelled as J-6P-0V, J-6P-45V and J-6P-90V in short form. $45\degree$ view and $90\degree$ view are rotated either clockwise or counterclockwise from the frontal view.
Appendix B Mathematical Formulations of Flow Snapshot Neurons
B.1 Slot Attention Module
As explained in Sec. 2.2, the slot attention module [63] aims to obtain flow snapshot neurons $\hat{Z}\in\mathbb{R}^{K\times D}$ to capture prototypical moments from the patch-level optical flow $\hat{O}\in\mathbb{R}^{S\times D}$ where $S=T\times N$ and $D=2\times(T-1)$. Formally, based on the cross-attention mechanism [98], slots $\hat{Z}$ serve as the query, $\hat{O}$ contribute to both the key and the value, and $q(\cdot),k(\cdot),v(\cdot)$ are the linear transformation employed to project inputs and slots into a common dimension $B$, which can be formulated as:
$\text{attn}_{i,j}:=\frac{e^{J_{i,j}}}{\sum_{l}^{K}e^{J_{i,l}}}\quad\text{ where }\quad J:=\frac{1}{\sqrt{D}}k(\hat{O})\cdot q(\hat{Z})^{T}\in\mathbb{R}^{S\times K},$ (3)
$\hat{Z}_{p+1}=\text{GRU}\left(\hat{Z}_{p},h_{p}\right).$ (5)
B.2 Contrastive Walk Loss
$L_{slot}=\text{CE}(Q_{(\hat{Z},\hat{O})}Q_{(\hat{O},\hat{Z})},\mathbf{I}),$ (6)
where $Q_{(a,b)}$ is defined as the normalized pairwise feature similarities between $a$ and $b$ in Sec. 2.1. We use the temperature $\mu$ of 0.05 in $Q_{(\hat{Z},\hat{O})}$ and $Q_{(\hat{O},\hat{Z})}$. CE($\cdot$,$\cdot$) stands for the cross-entropy loss and $\mathbf{I}\in\mathbb{R}^{K\times K}$ is the identity matrix. The effect of $\mu$ is covered in Appendix, Tab. S4.
Appendix C More Training Details
Our model is trained on the training sets of RGB videos in BMP, NTU RGBD+60 and NW-UCLA datasets respectively. BMP: Videos are resized to $224\times 398$ pixels for 50 epochs training with a batch size of 48. NTU RGB+D 60: Videos are resized to $224\times 398$ pixels and trained for 100 epochs with a batch size of 128. NW-UCLA: Videos are resized to $224\times 300$ pixels and trained for 100 epochs using a batch size of 16.
Figure S1: Schematic of human psychophysics experiments on Amazon Mechanical Turk (MTurk). In every trial, subjects are presented with a video and a list of ten options. After the video plays only once, the subject has to choose one action among ten options that best describe the action in the video.
Appendix D Human Psychophysics Experiments
Fig. S1 demonstrates the schematic of human psychophysics experiments on Amazon Mechanical Turk (MTurk)  [8]. We recruited a total of 90 human subjects. We collect data from a total of 12,600 trials.
In each trial, participants are shown a video randomly selected from the BMP dataset and are then asked to perform a forced 1-out-of-10 choice test to identify the action in the video. Videos from all conditions are randomly selected and presented in random order. The stimuli are uniformly distributed across BMP conditions, with no long-tailed distribution. These action classes are commonly performed in our daily life and free from cultural bias, as psychophysics experiments show humans can recognize them with nearly 100% accuracy on RGB videos. Almost all subjects are from the US, and we did not collect demographic data. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. Each subject was compensated.
To control data quality, two checks are implemented in the experiment: (1) Four pre-selected videos in the RGB condition are used as the dummy trials, with two videos representing the "Sit Down" action class and the other two representing the "Stand Up" action class. Since these four videos are meant for quality controls. There is no ambiguity in classifying the action classes in these four videos. These four dummy trials are randomly dispersed with the rest of the actual trials in one experiment. Participants who fail to recognize correct actions in any of the four videos are excluded for data analysis. 88 out of the total 90 participants passed the tests in the dummy trials. (2) Each participant is allowed to participate in the experiment only once. All the trials in one experiment are always unique. This is to prevent the subjects from memorizing the answers from the repeated trials.
Top-1 Acc. (%)
RGB J-6P SP-8P-1LT
E2S-X3D [42] 98.7 10.2 10.7
VideoMAE [94] 90.0 9.9 9.9
TwoStream-CNN [90] 97.0 15.7 10.4
MP(ours) 96.5 69.0 49.7
Table S1: Results of more baselines and our motion perceiver (MP) in BMP tasks. Our MP demonstrates superior performance than baselines on RGB videos, Joint videos with 6 points (J-6P), and SP videos with 8 points and a lifetime of 1 (SP-8P-1LT). Top-1 accuracy (%) is reported. Best is in bold.
# Trainable Parameters (M)
ResNet3D 31.7 SlowFast 33.7
I3D 27.2 X3D 3.0
R(2+1)D 27.3 MViT 36.1
MP(ours) 57.5
Table S2: Number of trainable parameters in million (M) for baselines and our MP. Although our model is larger than the baselines, its superior performance is not a result of the larger number of parameters.
Appendix E More Baselines Comparisons
Besides the comparison of State-of-the-Art methods discussed in Sec. 3.2, this section offers comparisons with three additional baseline methods. E2-S-X3D [42] is a two-stream architecture processing optical flow and spatial information from RGB frames separately. VideoMAE [94], a recent baseline trained on Kinetics in a self-supervised learning manner. TwoStream-CNN [90] incorporates the spatial networks trained from static frames and temporal networks learned from multi-frame optical flow. The results in Appendix, Tab. S2 indicate that our method significantly surpasses all baselines under the J-6P and SP-8P-1LT conditions, highlighting the superior generalization capability of our MP model on the BMP dataset.
Moreover, as stated in Sec. 3.2, we explore the upper bound performance of our MP on Joint and SP videos. First, we directly train our MP model on J-6P and test it on J-6P. Its accuracy on J-6P is 95% whereas human performance is 86%. Surprisingly, we found this model trained on J-6P also achieves 55% accuracy in RGB and 71% in SP-8P-1LT, which are far above chance. This implies that our model has generalization ability across multiple modalities. Second, we also train our model on SP-8P-1LT and test it on all three modalities: 43% in RGB, 69% in J-6P, and 93% in SP-8P-1LT. The reasonings and conclusions stay the same as the model directly trained on J-6P. Note that although our model achieves very high accuracy on the in-domain test set (train on J-6P, test on J-6P and train on SP-8P-1LT, test on SP-8P-1LT), its overall performance over all three modalities (RGB, J-6P, and SP-8P-1LT) is still lower than humans (74% vs 88%). This emphasises the importance of studying model generalisation in BMP. There is still a performance gap between AI models and humans in BMP.
Appendix F Enhanced Motion Perceiver (En-MP)
It is worth noting that our current MP model (see Sec. 2) is applied solely to the feature maps from the final attention block of DINO (block 12). We have observed that our MP model can also be effectively applied to early and intermediate-level blocks of DINO. The final prediction is then generated by fusing features across these three blocks (blocks 1, 7, and 12). We refer to this improved model as the Enhanced Motion Perceiver (En-MP).
Appendix G Viewpoint Discrepancy
As mentioned in Sec. 4.1, human behaviors in our study reveal a divergent trend compared to the human performance pattern described in [46]. We point out two reasons: first, the tasks differ between our study and [46]. Our study focuses on action recognition while the work [46] focuses on walking pattern discrimination. Second, we found the performance variations across different camera views depend on the action classes. As shown in Appendix, Fig. S2, humans exhibit the best performance when viewing “Arm Circles" actions from the frontal view, because the movement of the two arms will not overlap in the frontal view. For the videos from the action class “Point To Something", the 90°view is the best viewpoint since it results in longer movement trajectories of the arm than other viewpoints. Therefore, although humans can achieve the best average accuracy on the 45°viewpoint in our experiment, it does not imply that humans always perform the best on the 45°viewpoint in all action classes.
Figure S2: Human performance across camera views within different action classes on J-6P videos in BMP dataset. The labels on the x-axis are the 10 action classes in the BMP dataset (Appendix, Sec. A.2).
Appendix H Alignment Between Models and Humans
The alignment between models and humans can be assessed in three aspects: (1) The correlation coefficients between all the AI models and human performance across all the BMP conditions. The results in Fig. 6 show that our MP has the highest correlation with human performance compared with baselines. (2) The absolute accuracy in all action classes across all BMP properties, which is reported in Appendix, Fig. S3. We averaged the accuracies within the same property. It is observed that all AI models demonstrated lower accuracy than humans. However, among all AI models, our MP and En-MP outperform the rest. For example, despite explicitly modelling both motion and visual information in two separate streams, the SlowFast model still falls short in temporal orders. Similarly, MViT, which utilizes transformer architectures, fails to generalize to various temporal resolutions. See Appendix, Tab. S6 for detailed comparisons. (3) The error pattern consistency between AI models and humans using the metric introduced in [31]. The results in Appendix, Fig. S4 reveal that the error patterns from our MP and En-MP are more consistent with human error patterns than all the baselines at the trial level.
Figure S3: Our MP and En-MP significantly surpass all existing models on both Joint and SP videos. A correlation plot between model and human performances is provided, with accuracy averaged over conditions within each property on the same type of stimulus (Sec. 3.1). The black dash diagonal serves as a reference.
Figure S4: Our MP and En-MP show higher error consistency with human performance compared to all existing models. The Error Pattern Consistency [31] between each AI model and human performance across all BMP conditions is reported.
Appendix I Baselines with Optical Flow as Input
As mentioned in Sec. 4.2, the MP model explicitly uses patch-level optical flow to extract temporal information for classification, whereas the baselines do not. Thus, here we introduce augmented baselines where they are trained on pixel-level optical flows. Specifically, we use GMFlow [105, 106] to extract the pixel-level optical flow from RGB videos in the BMP dataset, and then convert them into three-channel videos based on the Middlebury color code [5]. Taking these optical flow videos in the training set as the input to baselines, the performances on optical flow videos in the testing set and J-6P are shown in Appendix, Tab. S3. It is evident that all baselines achieve high accuracy on optical flow videos in the testing set, indicating that they effectively learn to recognize actions from optical flow inputs. However, their best performances on J-6P are significantly worse compared to the MP model (30.83% vs 69.00% in Appendix, Tab. S6), suggesting that our MP learns more generalizable motion representations from patch-level optical flows.
[…]
Table S3: Results of baselines with optical flow as input. Top-1 accuracy is reported on optical flow videos (OF) and Joint videos with 6 points (J-6P). Best is in bold.
Appendix J Visualization of Patch-Level Optical Flow
As outlined in Sec. 4.2, we provide a visualization of patch-level optical flow in vector field plots across example video frames for “stand up” action in Appendix, Fig. S5. We can see that patch-level optical flow mostly happens in moving objects (the person performing the action) and captures high-level semantic features. Hence, they are more robust to perturbations in the pixel levels and more compute-efficient.
Figure S5: Visualization of patch-level optical flow in vector field plots across a sequence of example video frames from "stand up" action class in the BMP dataset. Optical flow from (a) Frame 1; (b) Frame 8; (c) Frame 16; (d) Frame 31 to Frame 32.
Appendix K More Ablation Studies